# An Effective Deployment of Diffusion LM for Data Augmentation in Low-Resource Sentiment Classification

### v1

@@ Abstract
Zhuowei Chen1, Lianxi Wang1, [^2] footnotemark: 2, Yuben Wu1, Xinfeng Liao1, Yujia Tian1, Junyang Zhong1 1School of Information Science and Technology, Guangdong University of Foreign Studies, Guangzhou, China. 2Guangzhou Key Laboratory of Multilingual Intelligent Processing, Guangzhou, China. wanglianxi@gdufs.edu.cn
@@ 1 Introduction
Diffusion- CLS (ours) Cor. [sad] Today, the [M] was a [M]. It was [M][M].
@@
To address these challenges, one effective approach is data augmentation (DA), which enriches the diversity of the dataset without explicitly collecting new data Feng et al. (2021). Classic rule-based DA methods employ logical modifications to obtain pseudo samples, such as EDA Wei and Zou (2019), and AEDA Karimi et al. (2021). Model-based DA methods develop rapidly as the transformer architecture dominates the NLP field, most of these methods execute DA through corrupt-then-reconstruct (CTR), as examples shown in Table 1. Namely, masked language model (MLM) Wu et al. (2019); Kumar et al. (2020), and GENIUS Guo et al. (2022) which applies BART as the sample generator. Also, Anaby-Tavor et al. (2020) finetunes GPT-2 and generates pseudo samples with label prompts, called LAMBADA.
@@
DiffusionCLS initially finetunes PLM with a diffusion objective, functioning as a sample generator, followed by training the TC model in a noise-resistant manner. By fine-tuning the diffusion LM, we can then input original samples with their crucial tokens corrupted and use the label as a generation prompt to get new samples. This method diversifies the original dataset by replacing strong label-related tokens and also steers the model towards producing high-quality pseudo samples that comply with the diversity-consistency rule.
@@ 3 Methodology
Figure 1 illustrates DiffusionCLS. The diffusion LM-based sample generator generates new samples for data augmentation, while the TC model is trained for the specific TC task. Label-Aware Prompting and Label-Aware Noise Schedule are crucial for training the sample generator, and Conditional Sample Generation and Noise-Resistant Training contribute to the training of the TC model.
@@ 4.1 Datasets and Baselines
Table 3: Performances of TC models on dataset SST-2 under the few-shot setting. $\dagger$ denotes that results are from previous research. All our results are collected from 5 runs with different seeds.
@@ 4.3 Results and Analysis
Results in Entire Data Mode. As shown in Table 2, in general, the proposed DiffusionCLS outperforms most DA methods on domain-specific dataset SMP2020-EWECT and India-COVID-X, especially under G/E policy. Notably, the DiffusionCLS positively impacts the TC model across all policies and datasets, which most baselines fail.
@@
Results in Partial Data Mode and Few-shot Settings. As shown in Figure 5 and Table 9 in Appendix C, the proposed DiffusionCLS method consistently improves the classification model. Notably, DiffusionCLS matches the PLM baseline performance on the Arabic SenWave dataset using only 50% of the data samples.
@@ A.1 Experiment Setup
The low-resource challenge in TC includes problems like insufficient annotated samples, domain-specific adaptation problems, and imbalanced distribution. To measure the capability of the proposed DiffusionCLS to mitigate these problems, we conducted experiments on three domain-specific datasets with respect to the problems mentioned above, as shown in Table 7.
@@ A.2 Implementation
For implementation, we take bert-base-uncased [^3] https://huggingface.co/google-bert/bert-base-uncased and chinese-roberta-wwm [^4] https://huggingface.co/hfl/chinese-roberta-wwm-ext from the huggingface platform respectively for English and Chinese dataset training.
Also, hyper-parameters settings of our work are demonstrated in Table 5 and Table 6.
@@ A.3 Datasets
Appendix B Baselines
  • Non-Generative Methods
  • Generative Methods
  • Representation Augmentation Methods
Appendix C Experiment Results with Partial Data Mode

### v2

@@ Abstract
Zhuowei Chen1, Lianxi Wang1,$2$Corresponding author., Yuben Wu1, Xinfeng Liao1, Yujia Tian1, Junyang Zhong1 1Guangdong University of Foreign Studies, Guangzhou, China. 2Guangzhou Key Laboratory of Multilingual Intelligent Processing, Guangzhou, China. wanglianxi@gdufs.edu.cn
@@ 1 Introduction
Diffusion- CLS (ours) Cor. [sad] Today, the [M] was [M][M]. It was [M][M].
@@
To address these challenges, one effective approach is data augmentation (DA), which enriches the diversity of the dataset without explicitly collecting new data Feng et al. (2021). Classic rule-based DA methods employ logical modifications to obtain pseudo samples, such as EDA Wei and Zou (2019), and AEDA Karimi et al. (2021). Model-based DA methods develop rapidly as the transformer architecture dominates the NLP field, most of these methods execute DA through corrupt-then-reconstruct (CTR), as examples shown in Table 1. Namely, masked language model (MLM) Wu et al. (2019); Kumar et al. (2020), and GENIUS Guo et al. (2022) which applies BART as the sample generator. Also, Anaby-Tavor et al. (2020) proposed LAMBADA, which finetunes GPT-2 and generates pseudo samples with label prompts.
@@
DiffusionCLS initially finetunes PLM with a diffusion objective, functioning as a sample generator, followed by training the TC model in a noise-resistant manner. By fine-tuning the diffusion LM, we can then input original samples with their crucial tokens corrupted and use the label as a generation prompt to get new samples. This method diversifies the original dataset by replacing strong label-related tokens and also steers the model towards producing high-quality pseudo samples that comply with the diversity-consistency rule. Also, experimental codes have been released on GitHub [^1] https://github.com/JohnnyChanV/DiffusionCLS.
@@ 3 Methodology
Figure 1 illustrates DiffusionCLS. The diffusion LM-based sample generator generates new samples for data augmentation, while the TC model is trained for the specific task. Label-Aware Prompting and Label-Aware Noise Schedule are crucial for training the sample generator, and Conditional Sample Generation and Noise-Resistant Training contribute to the training of the TC model.
@@ 4.1 Datasets and Baselines
Table 3: Performances of TC models on dataset SST-2 under the few-shot setting. $\dagger$ denotes that results are from previous research. All of our results are collected from 5 runs with different seeds.
@@ 4.3 Results and Analysis
Results under Entire Data Mode. As shown in Table 2, in general, the proposed DiffusionCLS outperforms most DA methods on domain-specific datasets SMP2020-EWECT and India-COVID-X, especially under G/E policy. Notably, the DiffusionCLS positively impacts the TC model across all policies and datasets, which most baselines fail.
@@
Results under Partial Data Mode and Few-shot Settings. As shown in Figure 5 and Table 9 in Appendix C, the proposed DiffusionCLS method consistently improves the classification model. Notably, DiffusionCLS matches the PLM baseline performance on the Arabic SenWave dataset using only 50% of the data samples.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Social Science Fund of China (No. 22BTQ045).
@@ A.1 Experiment Setup
The low-resource challenge in TC includes problems like insufficient annotated samples, domain-specific adaptation problems, and imbalanced distribution. To measure the capability of the proposed DiffusionCLS to mitigate these problems, we conduct experiments on three domain-specific datasets with respect to the problems mentioned above, as shown in Table 7.
@@ A.2 Implementation
For implementation, we take bert-base-uncased [^4] https://huggingface.co/google-bert/bert-base-uncased and chinese-roberta-wwm [^5] https://huggingface.co/hfl/chinese-roberta-wwm-ext from the huggingface platform respectively for English and Chinese dataset training. Also, hyper-parameters settings of our work are demonstrated in Table 5 and Table 6.
A.3 Datasets
Appendix B Baselines
  • Non-Generative Methods
  • Generative Methods
  • Representation Augmentation Methods
Appendix C Experiment Results with Partial Data Mode